# Phase retrieval on circles and lines

### v2

@@ Abstract.
Let $f$ and $g$ be analytic functions on the open unit disc ${\mathbb{D}}$ such that $|f|=|g|$ on a set $A$. We first prove that there exists $c$ in the unit circle ${\mathbb{T}}$ such that $f=cg$ when $A$ is the union of two lines in ${\mathbb{D}}$ intersecting at an angle that is an irrational multiple of $\pi$. The same conclusion is valid when $f$ and $g$ are in the Nevanlinna class and $A$ is the union of the unit circle and an interior circle, tangential or not. We also provide sequential versions of the previous results and analyse the case $A=r{\mathbb{T}}$. Finally we examine the situation when there is equality on two distinct circles in the disc, proving a result or counterexample for each possible configuration.
@@ 1. Background
In [7] Pohl, Li and Boche worked with the class $H^{1}_{R}$ of functions in $H^{1}=H^{1}({\mathbb{D}})$ such that the inner factor is a Blaschke product (no singular part). They showed (their Theorem 3) that any such function $f$ is determined to within a unimodular constant by the values of $|f|$ on ${\mathbb{T}}$ and $r{\mathbb{T}}$ for any $10<r<$ 1.
@@
This paper deals with several cases, for which it provides positive and negative results: intersecting curves, disjoint circles, disjoint lines, and a full analysis of the two-circle situation. We also show that many results have sequential counterparts.
@@ 2.1. Intersecting curves
The following result generalizes [3, Lem. 4.5], which proves the results for equality on two lines $(-1,1)$ and $e^{i\theta}(-1,1)$.
@@ 2.2. Disjoint circles
We can now prove a more general version of the main result, Theorem 3, from [7], removing the restriction on the singular factor.
@@
Let $f,g\in H^{1}({\mathbb{D}})$ satisfy $|f|=|g|$ a.e. on ${\mathbb{T}}$ and $r{\mathbb{T}}$ for some $10<r<$ 1. Then $f=\lambda g$ for some $\lambda\in{\mathbb{T}}$.
@@
The result above may easily be extended to the Nevanlinna class ${\mathcal{N}}$ of holomorphic functions expressible as the ratio of two $H^{1}$ functions.
@@
By Lemma 2.7 $|F|=1$ on ${\mathbb{T}}$ and so $|f|=|g|$ on $r{\mathbb{T}}$. The result now follows from Theorem 2.6.
@@ 2.3. Disjoint lines
Clearly, the inner factor $u$ has no zeros accumulating at a point on $i{\mathbb{R}}$. Moreover, the outer factor $h$ is continuous on the closed half-plane ${\mathbb{C}}_{+}$ (see, for example, the arguments in [6, Sec. 4.3.8]), which means that the inner factor $u$ is also continuous except possibly at the discrete set consisting of the zeros of $uh$ on $i{\mathbb{R}}$. On factoring out a zero at $iy_{0}$, by considering $uh/b$, where $b(s)=\dfrac{s-iy_{0}}{1+s}$ (an outer function) we see that $u$ is also continuous at the zeros of $uh$. Thus $u$ has no singular part except possibly an exponential factor $e^{-\alpha s}$ for some $\alpha\geq 0$. We conclude that $u(s)=B_{1}(s)e^{-\alpha s}$ where $B_{1}$ is a Blaschke product whose zeros do not accumulate at any point in $\overline{{\mathbb{C}}_{+}}$. The function $B_{1}$, which is a product of factors of the form $\dfrac{s-a_{n}}{s+\overline{a_{n}}}$, therefore has a meromorphic extension to ${\mathbb{C}}$.

### v3

@@ Abstract.
Let $f$ and $g$ be analytic functions on the open unit disc ${\mathbb{D}}$ such that $|f|=|g|$ on a set $A$. We give an alternative proof of the result of Perez that there exists $c$ in the unit circle ${\mathbb{T}}$ such that $f=cg$ when $A$ is the union of two lines in ${\mathbb{D}}$ intersecting at an angle that is an irrational multiple of $\pi$, and from this deduce a sequential generalization of the result. Similarly, the same conclusion is valid when $f$ and $g$ are in the Nevanlinna class and $A$ is the union of the unit circle and an interior circle, tangential or not. We also provide sequential versions of this result and analyse the case $A=r{\mathbb{T}}$. Finally, we examine the most general situation when there is equality on two distinct circles in the disc, proving a result or counterexample for each possible configuration.
@@ 1. Background
In [9] Pohl, Li and Boche worked with the class $H^{1}_{R}$ of functions in $H^{1}=H^{1}({\mathbb{D}})$ such that the inner factor is a Blaschke product (no singular part). They showed (their Theorem 3) that any such function $f$ is determined to within a unimodular constant by the values of $|f|$ on ${\mathbb{T}}$ and $r{\mathbb{T}}$ for any $10<r<$ 1. This restriction on the inner factor was removed by Perez [8].
@@
Finally, a recent paper of Liehr [5] looks at Gabor phase retrieval and Pauli-type uniqueness problems.
This paper deals with several cases, for which it provides positive and negative results: intersecting curves, disjoint circles, disjoint lines, and a full analysis of the two-circle situation. We also show that many results have sequential counterparts, thus extending and generalizing the work of [8].
@@ 2.1. Intersecting curves
Perez [8] proved a version of the following result, which generalizes [3, Lem. 4.5]. We give an independent proof, introducing a method that allows us to provide further extensions.
@@ 2.2. Disjoint circles
As described in the introduction to this paper, Perez [8] proved a more general version of the main result from [9], removing the restriction on the singular factor. We give an alternative proof, which leads us to more general results in the same area.
Recall that the Nevanlinna class ${\mathcal{N}}$ consists of those holomorphic functions in the unit disc ${\mathbb{D}}$ that are expressible as the ratio of two $H^{1}$ functions.
@@
Let $f,g\in{\mathcal{N}}$ satisfy $|f|=|g|$ a.e. on ${\mathbb{T}}$ and $r{\mathbb{T}}$ for some $10<r<$ 1. Then $f=\lambda g$ for some $\lambda\in{\mathbb{T}}$.
@@
By Lemma 2.7, $|F|=1$ on ${\mathbb{T}}$ and so $|f|=|g|$ on $r{\mathbb{T}}$. The result now follows from Theorem 2.6.
@@ 2.3. Disjoint lines
Clearly, the inner factor $u$ has no zeros accumulating at a point on $i{\mathbb{R}}$. Moreover, the outer factor $h$ is continuous on the closed half-plane ${\mathbb{C}}_{+}$ (see, for example, the arguments in [7, Sec. 4.3.8]), which means that the inner factor $u$ is also continuous except possibly at the discrete set consisting of the zeros of $uh$ on $i{\mathbb{R}}$. On factoring out a zero at $iy_{0}$, by considering $uh/b$, where $b(s)=\dfrac{s-iy_{0}}{1+s}$ (an outer function) we see that $u$ is also continuous at the zeros of $uh$. Thus $u$ has no singular part except possibly an exponential factor $e^{-\alpha s}$ for some $\alpha\geq 0$. We conclude that $u(s)=B_{1}(s)e^{-\alpha s}$ where $B_{1}$ is a Blaschke product whose zeros do not accumulate at any point in $\overline{{\mathbb{C}}_{+}}$. The function $B_{1}$, which is a product of factors of the form $\dfrac{s-a_{n}}{s+\overline{a_{n}}}$, therefore has a meromorphic extension to ${\mathbb{C}}$.